OpenStudy (anonymous):

please help! Are 2x+4 / 3x+6 and 2/3 equivalent expressions? Using complete sentences explain why or why not.

OpenStudy (raden):

factor out the numerator and the denominator (2x+4) / (3x+6) = 2(x+2)/3(x+2) cancel the x+2s = 2/3

OpenStudy (anonymous):

thank you!

OpenStudy (raden):

welcome

OpenStudy (anonymous):

so the (2x+4)/(3x+6) simplified is 2/3?

OpenStudy (raden):

yes

OpenStudy (anonymous):

okay thanks again :)